# AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence

### v2

@@
Abstract
Cyber Threat Intelligence (CTI) reports document observations of cyber threats, synthesizing evidence about adversaries’ actions and intent into actionable knowledge that informs detection, response, and defense planning. However, the unstructured and verbose nature of CTI reports poses significant challenges for security practitioners to manually extract and analyze such sequences. Although large language models (LLMs) exhibit promise in cybersecurity tasks such as entity extraction and knowledge graph construction, their understanding and reasoning capabilities towards behavioral sequences remains underexplored. To address this, we introduce AttackSeqBench, a benchmark designed to systematically evaluate LLMs’ reasoning abilities across the tactical, technical, and procedural dimensions of adversarial behaviors, while satisfying Extensibility, Reasoning Scalability, and Domain-dpecific Epistemic Expandability. We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across the proposed 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain. Our findings contribute to a deeper understanding of LLM-driven CTI report understanding and foster its application in cybersecurity operations.
1 Introduction
@@
As illustrated in Figure 1, we define the sequence of adversary behaviors as attack sequence (Al-Sada et al., 2025) to represent the execution flow of malicious actions across different stages of a cyber attack under the MITRE ATT&CK® framework (Strom et al., 2018). Building on this definition and the following key perspectives, we further delve into the suitability of LLMs in analyzing attack sequences. 1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate attack sequences from newly observed CTI reports. 2) Reasoning Scalability: Recently, Large Reasoning Models (LRMs) have demonstrated substantial advantages over conventional LLMs in multi-step reasoning tasks, such as coding and mathematical reasoning. However, existing CRU works have primarily focused on addressing CTI–related tasks via LLM, leaving the necessity of reasoning for attack sequence analysis largely unexplored. 3) Domain-Specific Epistemic Expandability: LLMs exhibited limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024b), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations. This requirement becomes particularly pronounced in attack sequence analysis, which necessitates a comprehensive understanding of adversarial behaviors to effectively reason multi-stage cyber attacks.
Figure 1: Illustration an example cyber attack sequence and our AttackSeqBench.
@@
Building upon these perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis. Catering to Extensibility, we first construct attack sequences based on extensive real-world CTI reports, ensuring that the benchmark accurately reflects the complexity and diversity of Tactics, Techniques, and Procedures (TTPs) in cyber attacks performed by APTs. Moreover, we design three Question Answering (Q&A) tasks under the adversary behaviors hierarchy in MITRE ATT&CK® and develop an automated Q&A generation pipeline that converts newly-collected CTI reports into the pre-defined format, enabling its extensibility on the corpus side. Following Reasoning Scalability, we further evaluate several LRMs and reasoning distillation strategies, which function well in general domains, to identify their strengths and limitations on the specialized attack sequence analysis task, providing helpful insights for future research in this area. To achieve Domain-Specific Epistemic Expandability, we aggregate cybersecurity-related knowledge from some existing benchmarks and embed it into LLMs via several post-training strategies to examine their epistemic expandability at the model level. Moreover, we also extend beyond the conventional zero-shot setting by introducing context-based and RAG-empowered settings, which pertinently assess LLMs’ epistemic expandability when injecting domain-specific cybersecurity knowledge at the semantic and representation levels.
@@
Our contribution are as follows: (I) We introduce AttackSeqBench, the pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to analyze attack sequences across diverse inference settings and multi-level tasks. (cf. Section 2) (II) We quantitatively demonstrate that existing LRMs fail to substantially outperform LLMs on attack sequence analysis and perform markedly worse in most cases, a contrast to their advantages observed in domains such as mathematics and coding. (cf. Section 3.3) (III) We offer a comprehensive analysis of how parameterization and parameter scale affect existing models’ attack sequence analysis, and further examine why current LRMs and RAG underperform on this specialized task. This work uncovers the fundamental limitations of current models in attack sequence analysis and provides actionable insights to guide future research in this domain. (cf. Section 3.4 and Section 3.5)
2 Dataset Construction and Verification
@@
2.1 Problem Definitions
CTI report understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge. To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB. Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
@@ 2.1 Problem Definitions
Figure 2: Overview of our automated QA dataset construction pipeline.
2.2 Dataset Construction
As illustrated in Figure 2, we first construct attack sequences using the extracted TTPs and CTI outline from CTI reports. Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the QA dataset.
Attack Sequence Construction. To efficiently and massively extract threat intelligence from these unstructured reports, we utilize a set of 408 CTI reports from various security vendors (Cisco Talos Intelligence Group, 2025; Microsoft, 2025) to construct attack sequences that accurately reflect the behaviors of real-world APTs. Specifically, we utilize a LLM-based KG construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combin them to construct the attack sequences $S$. Notably, we exclude CTI outlines which contains less than two ATT&CK tactics in attack sequence construction as they are unlikely to detail attack patterns observed in real-world cyber attacks.
@@ 2.2 Dataset Construction
For the Multiple-Choice Question (MCQ) tasks, we adopt a rule-based approach to select three choices as distractors. Specifically, we select a adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k+1}$ or $t_{k-1}$) and randomly select two tactics from the ATT&CK KB in AttackSeq-Tactic. Regarding AttackSeq-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first technique belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\notin E(t_{k})$; (2) The second technique is supported by the given attack sequence but belongs to another tactic $e_{j,j}\notin E(t_{k})$; (3) The third technique comes from a randomly chosen tactic that is not supported by the given attack sequence.
Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeq-Procedure-Yes. Next, we randomly sample its 70% questions to generate the negative question samples. Specifically, we design two types of Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” and/or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
@@
Considering the foundational role of Answerability in benchmark design, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives. Questions failing this requirement are discarded from the next step. Secondly, the LLM is instructed to evaluate the questions based on the remaining five aspects, providing a numerical score (out of five) and feedback for each aspect. Lastly, the LLM is prompted to refine the questions based on the feedback given (cf. Appendix C.2). We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final QA dataset.
@@
After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/1,393/3,249 to 1,697/1,917/1,223/1,412, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria. Additionally, we further illustrate the top-10 ATT&CK tactics and techniques within our dataset in Figure 7(a) and 7(b) respectively. The most frequent tactic and technique in the figure is associated with a key objective of APTs, highlighting the relevance of our Q&A dataset in capturing attack sequences based on real-world cyber attacks.
Table 1: Evaluation results of four sub-tasks on human evaluation and automatic evaluation.
Task Human Evaluation Automatic Evaluation
Hum. Perf. Ans. Cla. Log Rel. Con. Ans. Con. Ans. Cla. Log Rel. Con. Ans. Con.
Tactic 0.51 4.30 4.36 4.45 4.56 4.46 4.44 4.52 4.65 4.79 4.84 4.65 4.76
Technique 0.71 4.09 4.21 4.40 4.45 4.44 4.41 4.10 4.40 4.62 4.63 4.39 4.59
Procedure-Yes 0.74 4.88 4.70 4.88 5.00 4.81 4.94 4.02 4.06 4.61 4.47 3.78 3.89
Procedure-No 0.56 4.55 4.84 - - 4.82 4.66 3.29 3.66 - - 2.77 3.25
Average 0.63 4.45 4.53 4.57 4.67 4.63 4.61 3.98 4.19 4.67 4.64 3.90 4.12
2.3 Dataset Evaluation
@@
LLMs demonstrate strong potential in solving complex tasks, but they inevitably exhibit even severer hallucinations, which has become a widely recognized concern in the research community. To address this, we adopt a hybrid approach towards evaluating the quality of the constructed QA dataset using the criteria defined in our Q&A refinement (cf. Section 2.2). We design 5-point Likert scales for each of the evaluation criterion, where higher scores indicate better alignment.
@@
Human Evaluation. We first randomly sample 35 questions from each sub-task to construct a question set for human evaluation. Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 2.2. Based on Table 1, we observe that the average Human Performance (abbreviated as Hum. Perf.) equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise. Notably, AttackSeq-Procedure-No is derived from AttackSeq-Procedure-Yes through Yes-to-No transferring strategies, that is, its Logical and Relevance are inherently misaligned with the attack sequence, and we therefore do not evaluate these two aspects. Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.45$ to $4.61$ out of $5$, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
@@
Automatic Evaluation. To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain, including cybersecurity (Xu et al., 2024a). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated Q&A pairs. Specifically, we design individual prompts for each aspect in the evaluation criteria that includes its definition and the scoring guideline based on the same 5-point Likert scale in Human Evaluation (cf. Table LABEL:tab.annotation_instructions in Appendix A.2). Then we instruct GPT-4o rate the Q&A for each aspect based on the evaluation criterion and the correct answer from the ATT&CK KB (cf. Appendix C.3). Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences (cf. ). The fact that automatic evaluation scores are lower than human evaluation scores further indicates that answering questions correctly in our AttackSeqBench is more challenging for LLMs than for domain experts.
3 Benchmark and Experiments
Figure 3: Overview of the three benchmark settings that exhibit varying levels of contextual information given to the LLM.
3.1 Benchmark Settings

### v3

# AttackSeqBench: Benchmarking the Capabilities of LLMs for Attack Sequences Understanding

@@
Abstract.
Cyber Threat Intelligence (CTI) reports document observations of cyber threats, synthesizing evidence about adversaries’ actions and intent into actionable knowledge that informs detection, response, and defense planning. However, the unstructured and verbose nature of CTI reports poses significant challenges for security practitioners to manually extract and analyze such sequences. Although large language models (LLMs) exhibit promise in cybersecurity tasks such as entity extraction and knowledge graph construction, their understanding and reasoning capabilities towards behavioral sequences remains underexplored. To address this, we introduce AttackSeqBench, a benchmark designed to systematically evaluate LLMs’ reasoning abilities across the tactical, technical, and procedural dimensions of adversarial behaviors, while satisfying Extensibility, Reasoning Scalability, and Domain-dpecific Epistemic Expandability. We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain. Our findings contribute to a deeper understanding of LLM-driven CTI report understanding and foster its application in cybersecurity operations. Our code of benchmark construction and evaluation and the corresponding dataset are available at: https://github.com/hulkima/AttackSeqBench.
Cyber Security, Cyber Threat Intelligence, Large Language Models
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
1. Introduction
@@
Figure 1. Illustration of an example cyber attack sequence and our AttackSeqBench.
[…]
@@
Against this backdrop, we examine the suitability of LLMs for analyzing attack sequences from the following three perspectives. 1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate the attack sequences from newly observed CTI reports. 2) Reasoning Scalability: Recently, Large Reasoning Models (LRMs) have demonstrated substantial advantages over conventional LLMs in multi-step reasoning tasks, such as coding and mathematical reasoning. However, existing CRU works have primarily focused on addressing CTI–related tasks via LLMs, leaving the necessity of reasoning within LRMs for attack sequence analysis largely unexplored. 3) Domain-Specific Epistemic Expandability: LLMs have exhibited significant limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024c), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations. This requirement becomes particularly pronounced in attack sequence analysis, which necessitates a comprehensive understanding of adversarial behaviors to effectively reason multi-stage cyber attacks.
@@
Building upon these aforementioned perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis. Catering to Extensibility, we first construct attack sequences based on extensive real-world CTI reports, ensuring that the benchmark accurately reflects the complexity and diversity of Tactics, Techniques, and Procedures (TTPs) in cyber attacks performed by APTs. Moreover, we design three Question Answering (Q&A) tasks under the adversary behaviors hierarchy in MITRE ATT&CK® and develop an automated Q&A generation pipeline that converts newly-collected CTI reports into the pre-defined format, enabling its extensibility on the corpus side. Following Reasoning Scalability, we further evaluate several LRMs and reasoning distillation strategies, which function well in general domains, to identify their strengths and limitations on the specialized attack sequence analysis task, providing helpful insights for future research in this area. To achieve Domain-Specific Epistemic Expandability, we aggregate cybersecurity-related knowledge from some existing benchmarks and embed it into LLMs via several post-training strategies to examine their epistemic expandability at the model level. Moreover, we also extend beyond the conventional Zero-Shot setting by introducing Context-based and RAG-empowered settings, which pertinently assess LLMs’ epistemic expandability when injecting domain-specific cybersecurity knowledge at the semantic and representation levels.
@@
Our contributions can be summarized as follows: (I) We introduce AttackSeqBench, a pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to understand attack sequences across diverse settings and multi-level tasks. (cf. Section 3) (II) We quantitatively demonstrate that existing LRMs fail to substantially outperform LLMs on attack sequence analysis and perform markedly worse in most cases, a contrast to their advantages observed in domains such as mathematics and coding. (cf. Section 4.3) (III) We offer a comprehensive analysis of how parameterization and parameter scale affect existing models’ attack sequence analysis, and further examine why current LRMs and RAG underperform on this specialized task. This work uncovers the fundamental limitations of current models in attack sequence analysis and provides actionable insights to guide future research in this domain. (cf. Section 4.4 and Section 4.5)
2. Related Work
Automating CTI Report Understanding. With the increasing demands of cybersecurity operations and the breakthrough of LLMs, researchers have progressively explored their applicability within CRU (Zhang et al., 2024a). For instance, prior works have showcased the remarkable capabilities of LLMs in interpreting TTPs from the ATT&CK Knowledge Base (KB), surpassing the performance of some of the fine-tuned LMs with cybersecurity data (Fayyazi et al., 2024). Meanwhile, another line of work proposes LLM-driven threat intelligence Knowledge Graph construction frameworks, which utilize the threat-related entities and relations to describe CTI reports in a structural manner (Huang and Xiao, 2024; Cheng et al., 2024). However, the extent to which LLMs can understand and reason about the precise relations between adversary behavioral sequences described in CTI reports remains largely under-explored. In our work, we perform a holistic evaluation on various pre-trained LLMs, LRMs and fine-tuned LLMs in attack sequence analysis, from deducing high-level tactics to procedures described in CTI reports.
Benchmarking LLMs in Cybersecurity. Inspired by the remarkable open-world knowledge and complex inference ability within LLMs, various benchmarks have been proposed to evaluate its general capabilities in language understanding (Hendrycks et al., 2021b), math reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021). Regarding the cybersecurity domain, researchers start to benchmark the abilities of LLMs under such specialized setting, such as ethical hacking and compliance (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023). Targeting CRU-related tasks, SEvenLLM (Ji et al., 2024) explores LLMs’ abilities in threat-related entities extraction and summarizes reports from security vendors. SecBench (Jing et al., 2025) evaluates the knowledge retention and logical reasoning abilities of existing pre-trained LLMs from multiple languages and dimensions. Meanwhile, CTIBench (Alam et al., 2024) introduces five benchmark tasks to explore the threat entity attribution and cause-tracing abilities of LLMs within the security context.
Figure 2. Overview of our automated Q&A dataset construction pipeline.
3. Dataset Construction and Verification
3.1. Problem Definitions
Attack sequence understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge. To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB. Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
@@ 3.1. Problem Definitions
3.2. Dataset Construction
As illustrated in Figure 2, we first construct attack sequences using the extracted TTPs and CTI outline from CTI reports. Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the Q&A dataset of our AttackSeqBench.
Attack Sequence Construction. To efficiently and massively extract threat intelligence from these unstructured reports, we utilize a set of 408 CTI reports from various security vendors (Cisco Talos Intelligence Group, 2025; Microsoft, 2025) to construct attack sequences that accurately reflect the behaviors of real-world APTs. The gold standard for real-world attack sequences is elusive because CTI reports vary in quality and completeness and practitioners with different expertise workflows and threat modeling frameworks reconstruct sequences in inherently-divergent ways. Thjrerfore, we utilize a LLM-based Knowledge Graph (KG) construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combine them to construct the attack sequences $S$. Notably, we exclude CTI outlines which contains less than two ATT&CK tactics in attack sequence construction as they are unlikely to detail attack patterns observed in real-world cyber attacks.
@@ 3.2. Dataset Construction
For the Multiple-Choice Question (MCQ) tasks, we adopt a rule-based approach to select three choices as distractors. Specifically, we select an adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k-1}$ or $t_{k+1}$) and randomly select two tactics from the ATT&CK KB in AttackSeqBench-Tactic. Regarding AttackSeqBench-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first wrong option $e_{i,k}$ belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\in E(t_{k})$; (2) The second wrong option $e_{j,l}$ is supported by the attack sequence but belongs to another tactic $e_{j,l}\notin E(t_{k})$; (3) The third wrong option comes from a randomly chosen tactic that is not supported by the attack sequence.
Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeqBench-Procedure-Yes. Next, we randomly sample its 70% questions to generate the negative samples. Specifically, we design two Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
Table 1. Evaluation results of four sub-tasks on human evaluation and automatic evaluation. Here, Hum. Perf. = Human Performance, Ans. = Answerability, Con. = Consistency, Ans. Con. = Answer Consistency.
Task Human Evaluation Automatic Evaluation
Hum. Perf. Clarity Ans. Relevance Con. Ans. Con. Logic Clarity Ans. Relevance Con. Ans. Con. Logic
AttackSeqBench-Tactic 0.51 4.36 4.30 4.56 4.46 4.44 4.45 4.65 4.52 4.84 4.65 4.76 4.79
AttackSeqBench-Technique 0.71 4.21 4.09 4.45 4.44 4.41 4.40 4.40 4.10 4.63 4.39 4.59 4.62
AttackSeqBench-Procedure 0.64 4.78 4.70 - 4.82 4.79 - 3.85 3.63 - 3.24 3.55 -
Average 0.63 4.51 4.43 4.50 4.63 4.60 4.42 4.19 3.97 4.72 3.90 4.13 4.69
@@
Considering the foundational role of Answerability, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives. Questions failing this requirement are discarded from the next step. Secondly, the LLM is instructed to evaluate the questions based on the remaining five aspects, providing a numerical score (out of five) and feedback for each aspect. Lastly, the LLM is prompted to refine the questions based on the feedback given (cf. Appendix C.2). We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final Q&A dataset.
@@
After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/4,642 to 1,697/1,917/2,635, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria. Additionally, we further illustrate the top-10 ATT&CK tactics and techniques within our dataset in Figure 7 (a) and 7 (b) respectively. The most frequent tactic and technique in the figure is associated with a key objective of APTs, highlighting the relevance of our Q&A dataset in capturing attack sequences based on real-world cyber attacks.
$3.$ Dataset Evaluation
@@
LLMs show potential in solving complex tasks, but they inevitably exhibit severer hallucinations, which has become a widely recognized concern in the research community. To address this, we adopt a hybrid approach towards evaluating the quality of the constructed Q&A dataset using the criteria defined in our Q&A refinement (cf. Section 1). We design 5-point Likert scales for each of the evaluation criterion, where higher scores indicate better alignment.
@@
Human Evaluation. We first randomly sample 35 questions from each sub-task to construct a question set for human evaluation. Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 1. Based on Table 1, we observe that the average Human Performance equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise. Notably, AttackSeqBench-Procedure-No in AttackSeqBench-Procedure is derived from AttackSeqBench-Procedure-Yes via Yes-to-No transferring strategies, where its Logical and Relevance are inherently deviate from the given attack sequence, and we therefore do not evaluate these two aspects. Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.43$ to $4.63$ out of 55, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
@@
Automatic Evaluation. To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain (Xu et al., 2024b). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated Q&A pairs. Specifically, we design individual prompts for each aspect in the evaluation criteria that includes its definition and the scoring guideline based on the same 5-point Likert scale in Human Evaluation (cf. Table 5 in Appendix A.2). Then we instruct GPT-4o rate the Q&A for each aspect based on the evaluation criterion and the correct answer from the ATT&CK KB (cf. Appendix C.3). Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences. The fact that automatic evaluation scores are lower than human evaluation scores further indicates that answering questions correctly in our AttackSeqBench is more challenging for LLMs than for domain experts.
4. Benchmark and Experiments
4.1. Benchmark Settings